OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9> help please!!

jimthompson5910 (jim_thompson5910):

use the formula \[\Large \theta = arcos\left( \frac{u \cdot v}{|u||v|}\right )\]

jimthompson5910 (jim_thompson5910):

where |u| = length of vector u |v| = length of vector v

OpenStudy (anonymous):

I got 71.6degrees is that correct?

jimthompson5910 (jim_thompson5910):

close enough, I'm getting 71.56505117707799 which rounds to 71.6

OpenStudy (anonymous):

okay thank you!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

could you help me with this one too? Given that P = (-7, 16) and Q = (-8, 7), find the component form and magnitude of QP

jimthompson5910 (jim_thompson5910):

if you want QP, then you subtract the corresponding coordinates x: -7 - (-8) = -7 + 8 = 1 y: 16 - 7 = 9 So the vector QP is <1,9> If you wanted vector PQ, you would flip the signs to get <-1,-9> since it's just going in the complete opposite direction

jimthompson5910 (jim_thompson5910):

The magnitude of QP is equal to sqrt(x^2+y^2) where x and y are the components of the vector

OpenStudy (anonymous):

okay I got.... (1, 9), 82

jimthompson5910 (jim_thompson5910):

should be \[\Large \sqrt{82}\]

OpenStudy (anonymous):

okay, thank you! on this question I keep getting 28, -6....but i know thats wrong.....Let u = <-6, 3>, v = <-1, -6>. Find 4u + 2v.

jimthompson5910 (jim_thompson5910):

that's not correct

jimthompson5910 (jim_thompson5910):

4u means you multiply each component of vector u by 4

jimthompson5910 (jim_thompson5910):

u = <-6, 3> 4u = 4*<-6, 3> 4u = <4*(-6), 4*3> 4u = <-24, 12>

jimthompson5910 (jim_thompson5910):

Do the same for 2v

jimthompson5910 (jim_thompson5910):

Then add the resulting vectors

OpenStudy (anonymous):

-26,0 got it...still up for more help :) lol

jimthompson5910 (jim_thompson5910):

<-26, 0> is correct

jimthompson5910 (jim_thompson5910):

and I can do one more

OpenStudy (anonymous):

Determine whether a triangle can be formed with the given side lengths. If so, use Heron’s formula to find the area of the triangle. a = 240 b = 121 c = 302

jimthompson5910 (jim_thompson5910):

if a triangle with side lengths a, b,c could be formed, then ALL of the following must be true a+b > c a+c > b b+c > a

jimthompson5910 (jim_thompson5910):

basically, the sum of any two sides must be larger than the third side

OpenStudy (anonymous):

so no triangle can be form?

OpenStudy (anonymous):

nevermind..

OpenStudy (anonymous):

so what's next

jimthompson5910 (jim_thompson5910):

let's see a+b > c 240+121 > 302 361 > 302 So that's true

jimthompson5910 (jim_thompson5910):

a+c > b 240+302 > 121 542 > 121 that's true as well

jimthompson5910 (jim_thompson5910):

b+c > a 121+302 > 240 423 > 240 and that's true

jimthompson5910 (jim_thompson5910):

So all 3 are true, meaning that a triangle could be formed

OpenStudy (anonymous):

now how do I got about finding the area..

jimthompson5910 (jim_thompson5910):

Now you must find the semi-perimeter S S = (a+b+c)/2

OpenStudy (anonymous):

332.5

jimthompson5910 (jim_thompson5910):

now use that, along with a,b,c in the formula below A = sqrt( S*(S-a)*(S-b)*(S-c) ) this is heron's formula

jimthompson5910 (jim_thompson5910):

and it's used to find the area of any triangle if you know all 3 sides (it doesn't have to be a right triangle and you don't have to find the height)

OpenStudy (anonymous):

24,172.48 correct?

jimthompson5910 (jim_thompson5910):

let me check oh and btw, you can read more about heron's formula here http://en.wikipedia.org/wiki/Heron%27s_formula

jimthompson5910 (jim_thompson5910):

no that's way too high

OpenStudy (anonymous):

ahhh I see what I did... 13,724.27?

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

thank you sooooo much for all the help! I can't thank you enough!!

jimthompson5910 (jim_thompson5910):

yw